OpenStudy (anonymous):

What is the interest on $250.00 at 4% for three yea rs?

OpenStudy (anonymous):

simple or compound?

OpenStudy (anonymous):

Nope. Unfortunately.

OpenStudy (anonymous):

Doesn't say?

OpenStudy (anonymous):

by simple interest it would be\[Interest = $250 \times 0.04 \times 3\]

OpenStudy (anonymous):

I got $281.21. You earned $31.21 This is what I got

OpenStudy (anonymous):

I think it is Amount = 250 (1 + 0.04)^3 Because the way you put it doesnt make sense cause I dont get a logical answer

OpenStudy (anonymous):

yeah u r right it would be \[Interest = ($ 250 \times (1.04)^3) - $ 250\]

OpenStudy (anonymous):

\[250(.104)^3=281.21\] rounded so the interest is $32.21

OpenStudy (anonymous):

lol i guess i meant $31.21

OpenStudy (anonymous):

Thank you. lol (:

OpenStudy (anonymous):

yw